# Information Access of the Oppressed: A Problem-Posing Framework for Envisioning Emancipatory Information Access Platforms

### v1

@@ 3. Information Access of the Oppressed
This designation of information seekers as agency-free “objects” normalizes the platform monitoring their mouse clicks (Chuklin et al., 2015) (and even mouse movements (Diaz et al., 2013)) to farm signals and training data for relevance models. The reduction of individual users to their interactions with the system incentivizes their manipulation even by non-platform actors—e.g., the growing practices of clickbaiting (O’Donovan, 2014) and rage-baiting (Hom, 2015) by content producers. Furthermore, even in the face of serious societal concerns this frame restricts our imagination to only be able to conceive the same technologists—or a subset of their altruistic converts—as those that must bear the burden of mitigating the risks to society and its marginalized populations. Seeing through this frame one comes to the inevitable (but false) conclusion that all mitigative research must then focus solely on inventing mechanisms that IA platforms can operationalize to address concerns of bias, fairness, misinformation and disinformation, privacy, and safety; and we must ensure platforms employ said mechanisms by appealing to their better nature, through public pressure, or via enforcement of legal regulations. But what remains elusive just beyond the reach of this frame is the idea that may be the control over these platforms should never have been centralized in the hands of large corporations, because that threatens Big Tech’s grand project of empire building (Hao, 2025). To those well-intentioned “converts”, we quote the following text by Freire:
@@
This is not an argument against research (or researchers) that develop mitigative intervention mechanisms, but rather a call for us to recognize that those efforts must be grounded in broader collective acts of movement-building to challenge the concentration of power in the hands of large (often multinational) platforms and to ultimately abolish Big Tech. Otherwise, any positive impact of these efforts will be temporary and potentially contribute to ethics washing (Wagner, 2018; Van Maanen, 2022) and further entrenchment of corporate hold over our IA platforms. Instead, we advocate that societally-motivated technologists should join the struggle in solidarity with the people who are most at risk and harmed by these technologies. With this re-framing, the role of the technologists becomes two-fold. They must present the material risks of emerging technologies for the marginalized as an object of study to the people (users), to encourage them to see themselves as the necessary makers and unmakers of technology in support of their liberation. Simultaneously, technologists must re-envision the structure of our online technology stacks, such that they facilitate marginalized communities co-opting and co-constructing the technology to build pedagogical information experiences that encourage information seekers to critically examine their relations with oppressive structures, and to take action in support of universal emancipation. The inclusion of those at the cross hairs of technology is therefore much more than symbolic. It is only through their participation that we can aim for the necessary social, political, and cultural sensitivity in our pursuit for universal emancipation. Their participation is a radical act of connecting their struggles with the making and unmaking of technology, that must be shaped by intra-community and inter-community negotiations. This co-construction of technology is both informed by social movements and itself becomes an integral part of liberatory movement-building.
@@
When designing new modes of participation to center the needs of the marginalized, we must counter cynicism about the likelihood of broad public engagement by reminding ourselves that communities have long been resisting and appropriating everyday technologies to alleviate their oppression (Nemer, 2022). It is true that the ability to appropriate and rework technology is considerably more limited in the age of online services, since the “code” and infrastructure is not easily accessible to or modifiable by users, relative to the ease of appropriating physical devices. Therefore, we cannot assume a reluctance to engage, but must rethink online technology stacks to re-expose those cracks, transforming them into legitimate spaces for not just participation, but appropriation in support of everyday struggles against oppression. These spaces must be structurally guaranteed in our design to support people marginalized by these technologies to reflect, negotiate, and collaborate with each other to co-opt and co-construct the technology to suit their own needs.
@@
To re-construct technology in support of emancipation, we must work together to unequivocally reject the premise of technological determinism (Chandler, 1995) and confront the repeating cycles of technodeterministic hype coming from our own technology communities—e.g., just in recent years: Blockchain, Non-fungible tokens (NFTs), Virtual Reality, and now AI. This is not to deny that many of these technologies have their material usefulness, [^12] OK, may be not NFTs. (Yang, 2023) but the ask is to recognize that each of them drove a frenzied hype that tried to convince the world that the ubiquity of the technology was inevitable. Consequently, we must also break the ever-repeating cycles of critiquing emerging technologies and developing incomplete means of mitigation when our resistance is better directed at those who wield power over these technologies to serve their own need for wealth and power. Instead of only obsessing about mitigations, we must build for something better, something for our collective selves, something hopeful and emancipatory. To quote Freire:
@@
So, to operationalize Freire’s principles we adopt a problem-posing approach to rethink IA platforms. As the instigators of this work—and as recovering “converts” ourselves who previously approached the sociotechnical concerns of IA through the traditional lenses of fairness, transparency, and related frames—we see our new role as that of problem-posing, to build collective movements for universal emancipation, to free our IA platforms from authoritarian control, and in turn to liberate ourselves from the complicity in the global rise in authoritarianism, following Freire’s guidance:
Problem-posing education, as a humanist and liberating praxis, posits as fundamental that the people subjected to domination must fight for their emancipation.
4. Problem-Posing Framework for Emancipatory IA
@@
Our framework comprises of a series of problem-posing questions selected to prompt participants to critically reflect, ideate, prototype, experiment, and study different types of emancipatory IA approaches and capabilities. The problems are chosen to reflect some of the primary concerns of emancipatory IA while minimizing assumptions and prescriptions about how these problems are to be addressed. We intend for each of these problems to solicit a diverse range of candidate sociotechnical design ideas and interventions that foregrounds our emancipatory aspirations. Each problem may be considered in isolation, but ultimately we seek visions for IA platforms that can jointly address as many of these concerns as possible. We aim for this framework of problems to expose a rich design space of alternative approaches that communities can explore to challenge the status quo IA platforms. Next, we enumerate our framework of questions. We group these questions into two thematic groups, focusing on (1) the operations and governance of an envisioned emancipatory IA platform, and (2) safeguarding such a platform from capture and co-option.
@@
Theme: Platform operation and governance. We pose a provocation to IA researchers and platform developers to rethink the design of our technological stacks to create the conditions for community co-option and co-construction of IA technologies. In addition, the platform must also explore models of democratic governance and effective mechanisms of policy enforcements. We also need research and prototyping—informed by multiple disciplines including HCI, information sciences, and critical theory—to explore the design space of pedagogical information experiences. To summarize, this theme concerns with models for community co-option and co-construction, democratic governance, policy enforcement, and design of pedagogical information experiences.
@@
    How do we design and develop the platform from the margins by legitimizing community co-option and co-construction of the platform?
    Desiderata. The platform must legitimize its own co-option and co-construction through reflective participation of community members. Community members should be empowered to build new IA artifacts (e.g., information retrieval components, moderation services, and information experiences). Community members should be able to leverage artifacts built by other communities, and may also build competing artifacts if existing ones are inadequate or deemed harmful to community members. Through these participatory interactions the platform must allow for intra- and inter-community negotiations. The platform must mediate these interactions to ensure an equal playing-field for community members. Participation in co-construction must not be biased by who has access to more resources and labor. The model of engaging communities must be informed by theories of power and strive for epistemic justice. Finally, to ensure the platform does not simply extract free labor from already marginalized communities, it must also consider mechanisms for their compensation and long-term sustainability of this work.
@@
    Desiderata. Community participation is not a miracle-cure for all our sociotechnical challenges with IA. The opportunity to co-opt and co-construct may not only be leveraged by the oppressed but also the oppressor. We see these tensions already in context of app stores where apps that enable racial surveillance are allowed to exist (Molla, 2019) but those that oppose authoritarian violence are pulled down under government pressure (Allyn, 2025). We must recognize the limits of community negotiations and need for platform governance to bend the arc of platform development towards emancipation. All platforms must also moderate (Gillespie, 2018) and platform governance must deal with critical moderation policies and mediate challenging cases for moderation. Content moderation and governance are not at odds with emancipatory aspirations; mitigating real-world harms is a core tenet of anti-oppression and emancipatory praxis. The challenge is to develop a critical democratic framework for decision-making with respect to platform governance that is also robust to capture. Any appropriate governance structures must ensure effective and timely removal of critically harmful content from circulation—e.g., child sexual abuse material (CSAM)—as well as to exercise sociopolitical sensitivity in decision-making for challenging cases like the sharing of the ‘Napalm girl’ photo (Levin et al., 2016).
@@
    Desiderata. Democratic decision-making must be supported by robust enforcement of its governing policies and decisions. For critical cases, like that of CSAM, the model for empowering communities to co-construct the platform must not be a loophole to circumvent moderation. However, there is a natural tension here between developing infrastructure for the enforcement of policies and recreating infrastructure of censorship and centralized ideological control. To address this tension, we need innovations in sociotechnical infrastructure; and radical proposals for harm mitigation and justice that are restorative, reparative, and transformative (Kaba, 2021).
@@
    How do we design information experiences to promote emancipatory pedagogy?
    Desiderata. Emancipatory IA poses the challenge to imagine new ways for us to experience information that are pedagogical and foreground critical scholarship. Pedagogy in this context must not replicate the “banking concept of education”—nor act as tools for indoctrination—but encourage reflection, critical thinking, and dialog. This work must be informed by multidisciplinary perspectives including HCI, design, critical theory, social and political sciences, humanities, and arts.
Theme: Safeguards from capture and co-option. No struggle for liberation has the luxury to step outside of its own cage to architect a new liberated reality; the work begins from within the bounds of the oppressive system to transform the state of oppression to one of emancipation. The work on emancipatory IA must likewise begin situated within present-day structures of oppression. This requires our work to prioritize safeguarding the platform from capture, blocking, political pressure, economic competition, and co-option. We consider these problems under this theme.
@@
    Desiderata. The platform must ensure that it cannot be pressured to take an action (or be blocked from taking an action) via state regulations or orders, through corporate or other institutional mandates, using economic and other incentives, by threats and acts of force or violence, and other forms of coercion. This requires that the infrastructure and its management, the governance of the platform, its sources of funding, and other critical aspects be decentralized to avoid exposing critical points of weakness that oppressive-actors can put pressure on to capture the platform. Any serious attempt at building an emancipatory IA platform must strategically consider adversarial strategies of platform capture and develop necessary mitigation.
@@
    Desiderata. The platform should support reasonably convenient mechanisms to circumvent attempts by governments or Internet service providers to block access to the service. There are precedents for online platforms—e.g., the messaging app Signal (Whittaker, 2024)—that provide mechanisms for circumventing authoritarian block on the app. Emancipatory IA must support similar mechanisms to ensure continued access to information under authoritarian censorship.
@@
    Desiderata. Any platform that succeeds in, or even attempts to, aid liberatory struggles would likely come under immediate political pressure. Beyond safeguarding against capture, the project of emancipatory IA must also consider how to ensure the physical and economic safety of those conducting the research and actively participating in development of the platform. This includes the safety of community members who participate in the co-construction of the platform. Safeguards may range from funding support, internal policies, community norms, technological safeguards, and other mechanisms. The platform would also face economic competition from other platforms, and capitalist corporations might be particularly motivated make the economics of the emancipatory IA platform untenable for its survival. Therefore, the economics of the platform should be a critical focus of the project. The platform strategy must adopt robust strategies to overcome these external socioeconomic and political forces to be sustainable for the long-term.
@@
    Desiderata. If the platform were to succeed, an immediate concern would be that it might get co-opted by authoritarian and corporate actors. Other existing IA platforms may also attempt to extract economic value from the emancipatory IA platform without reciprocating. The platform design must be intentional in safeguarding against these risks and codify the need for reciprocity through technological design, and licensing and other legal means—e.g., Copyleft licensing (Frantsvog, 2012). Emancipatory IA should draw insights from the successes and failures of past and current initiatives like the Open Source movement.
@@ 5. The Salient Tensions of Emancipatory IA
Emancipatory IA situates itself within the tension between the oppressor and oppressed; and resolves that tension by making a normative commitment to side with the oppressed. However, attempting to simultaneously resolve several of the problems posed in Section 4 evokes other critical tensions—spanning the technological, social, political, economic, epistemic, legal, and personal. These tensions will manifest as we collectively put the theories to practice. In this section, we briefly discuss a short selection of such tensions that we must negotiate.
@@
In response to the growing concentration of power of a handful of platforms over public speech and information access, there is an emerging debate over the platform vs. protocol approach (Masnick, 2019). This is particularly dominant in the social media discourse where protocols like ActivityPub and ATProto have gained popularity as alternatives to billionaire-owned social media platforms. In this paper, we have liberally used the term “platform” to refer to the object of our envisioning in the context of emancipatory IA. This is intentional, partly to sidestep this debate earlier in the paper and use “platform” as a shorthand to refer to whatever solution that emerges as appropriate but also to reflect the realities of building for users who may not understand or care about the mechanics of platforms vs protocols, as we discuss below. While protocols present distinct advantages over platforms for user ownership of data, competition, and resistance to single points of capture, as Hendrix and Flowers (2022) argue, protocols in lieu of platforms do not inherently challenge the hierarchical social structures that they emerge from; the benefits that protocols do provide often rest on a neoliberal belief that market competition will drive better outcomes. However, Emancipatory IA nudges us to consider democratic governance designed explicitly to drive towards emancipatory outcomes.
@@
Platform governance (Gorwa, 2019) is a critical challenge for online platforms. Emancipatory IA not only inherits these challenges but must also be hardened against the risk of capture of any governance mechanism by authoritarian actors. It therefore requires additional safeguards and safety protocols, as well as contingency plans for when members of the governing body may be targeted or coerced. Policy enforcement mechanisms must navigate a similarly challenging landscape, balancing effectiveness of enforcement and the risks of implementing the tools for its own capture. This is one of the core tensions that future research on emancipatory IA must resolve.
@@
Design-related tensions emerge from opening up the platform to community co-option and co-construction. When large masses of Twitter users abandoned the social media site in 2022 to flock to Mastodon, they experienced significant growing pains onboarding onto the new platform (Diaz, 2022). New users had to shoulder the burden to familiarize themselves with a new platform that posed significant friction, especially for non-tech-savvy users, and exercise trust-based actions (such as selecting a Mastodon instance) that many felt ill-equipped to do (Stefanija, 2023). In enabling, co-option and co-construction, emancipatory IA must still ensure not to offload responsibilities to the user that might overwhelm them, and continue to prioritize the ease of usability of the platform. It must also avoid fragmentation and drive towards consistency of user experience. Fundamentally, it must drive a balance between the flexibility of design it affords to contributing communities and the overall usability of the platform.
@@
Lastly, we must embark on this mission clear-eyed about challenges that emerge when the law diverges from justice. We opened this article with examples of authoritarian attempts at capture using executive orders and mandates. Under the conditions of oppression, the law disproportionately sides with the oppressor, which may create situations where the mandate of Emancipatory IA conflict with existing laws. Emancipatory IA platforms must find a robust protocol for decision-making to navigate these tensions while respecting legal frameworks where possible, human rights, and our social commitment to do no harm.

### v2

@@ 3.1. The technologist-user dichotomy
This designation of information seekers as agency-free “objects” normalizes platforms monitoring their online activities, including their clicks (Chuklin et al., 2015) and mouse movements (Diaz et al., 2013) to farm signals for evaluation and training data. The reduction of individuals to their interactions with the system incentivizes their manipulation even by non-platform actors—e.g., the growing practices of clickbaiting (O’Donovan, 2014) and rage-baiting (Hom, 2015) by content producers. Furthermore, this frame restricts our imagination to only be able to conceive the same technologists—or a subset of their altruistic converts—as those that should bear the burden of mitigating the risks to society and its marginalized populations.
3.2. Challenging the technologists-as-liberator frame
Our dominant imaginaries of technology development frame the technologists as the sole executors of technological change. Seeing through this frame one comes to the inevitable (but incorrect) conclusion that all research aiming to mitigate societal risks should focus solely on inventing mechanisms that IA platforms can operationalize to address concerns of bias, fairness, misinformation and disinformation, privacy, and safety; and we should ensure platforms employ said mechanisms by appealing to their better nature, through public pressure, or via enforcement of legal regulations. But what remains elusive just beyond the reach of this frame is the idea that may be the control over these platforms should never have been centralized in the hands of large corporations, because that threatens Big Tech’s grand project of empire building (Hao, 2025). To those well-intentioned “converts”, we quote the following text by Freire:
@@ 3.2. Challenging the technologists-as-liberator frame
This is not an argument against research (or researchers) that develop mitigative intervention mechanisms, but rather a call for us to recognize that those efforts should be grounded in broader collective acts of movement-building to challenge the concentration of power in the hands of large (often multinational) platforms and to ultimately abolish Big Tech. Otherwise, any positive impact of these efforts will be temporary and potentially contribute to ethics washing (Wagner, 2018; Van Maanen, 2022) and further entrenchment of corporate hold over our IA platforms. Instead, we advocate that societally-motivated technologists should join the struggle in solidarity with the people who are most at risk and harmed by these technologies. Dismantling the technologists-as-liberator frame will further require radical rethinking of our approach to technology design, and we operationalize Freire’s approach next to put forward such a proposal.
3.3. Freirean design
We define Freirean design (FD) as a design methodology and a political commitment to engage users in the active co-option and co-construction of technologies. The inclusion of those at the cross hairs of technology is much more than symbolic. It is an act that recognizes that epistemic and representational justice cannot be “gifted” by technologists to marginalized communities. It is also driven by the realization that the historical and cultural context of marginalization and the structures of oppression are not universal—and differ radically, while sharing some commonalities, from the inner city neighborhoods in the US to the favelas of Rio to the bastis of Kolkata. It is only through community participation that we can truly aim for the necessary social, political, and cultural sensitivity in our pursuit for universal emancipation. We should also be pragmatic that technologists—many of who are employed in the for-profit technology industry or depend on research funding from the military-industry complex—are unlikely to take the necessary risks to execute revolutionary change to challenge structural oppression faced by others‘out there’. But for community members, their participation becomes a radical act of connecting their existential struggles against oppression with the making and unmaking of technology, and in turn their participation becomes a site for solidarity and movement building.
@@ 3.3. Freirean design
When designing new modes of participation to center the needs of the marginalized, we should counter cynicism about the likelihood of broad public engagement by reminding ourselves that communities have long been resisting oppression and appropriating everyday technologies to alleviate their oppression (Nemer, 2022). It is true that the ability to appropriate and rework technology is considerably more limited in the age of online services, since the “code” and infrastructure is not easily accessible to or modifiable by users, relative to the ease of appropriating physical devices. Therefore, we cannot assume a reluctance to engage, but should refocus our energies in identifying structural barriers to their participation and rethink online technology stacks to re-expose the cracks that can be transformed into legitimate spaces for not just participation, but appropriation in support of everyday struggles against oppression. These spaces should be structurally guaranteed in our design to support people marginalized by these technologies to reflect, negotiate, and collaborate with each other to co-opt and co-construct the technology to suit their own needs.
@@
To re-construct technology in support of emancipation, we should also work together to unequivocally reject the premise of technological determinism (Chandler, 1995) and confront the repeating cycles of technodeterministic hype coming from our own technology communities—e.g., just in recent years: Blockchain, Non-fungible tokens (NFTs), Virtual Reality, and now AI. This is not to deny that many of these technologies have their material usefulness, [^15] OK, may be not NFTs. (Yang, 2023) but the ask is to recognize that each of them drove a frenzied hype that tried to convince the world that the ubiquity of the technology was inevitable. Consequently, we should also break the ever-repeating cycles of critiquing emerging technologies and developing incomplete means of mitigation when our resistance is better directed at those who wield power over these technologies to serve their own need for wealth and power. Instead of only obsessing about mitigations, we should build for something better, something for our collective selves, something hopeful and emancipatory. To quote Freire:
@@
So, in this new frame, the role of the technologist becomes two-fold. They should present the material risks of emerging technologies for the marginalized as an object of study to the people (users), mirroring Freire’s problem-posing approach, to encourage them to see themselves as the necessary makers and unmakers of technology in support of their liberation. Simultaneously, technologists should re-envision the structure of our online technology stacks—expanding on the ethos of ‘underdesigning’ systems at design time and putting critical emphasis on ‘designing the design process’ as popularized by proponents of the metadesign movement—such that they facilitate marginalized communities co-opting and co-constructing the technology to build pedagogical information experiences that encourage information seekers to critically examine their relations with oppressive structures, and to take action in support of their emancipation. As the instigators of this work—and as recovering “converts” ourselves who previously approached the sociotechnical concerns of IA through the traditional lenses of fairness, transparency, and related frames—we see ourselves in solidarity with the marginalized engaged together in the struggle to free our IA platforms from authoritarian control and to rid ourselves from our complicity in the global rise in authoritarianism. And in our new role, we recognize the importance of the practices of problem-posing in the process of affecting social and technological transformation.
4. Problem-Posing for Emancipatory IA
Applying FD to IA requires two types of design deliberations, although the boundaries between them may not always be distinct. First, the deliberation should focus on the overall platform design whose concerns include safeguarding against capture, strategic underdesigning of the platform to make space for community co-construction, and developing democratic frameworks for governance. The second type of deliberation should focus on information experience design aimed at epistemic and representational justice for intersectional communities—e.g., information experiences for teaching queer Dalit history. We recommend that both types of deliberations should operationalize Freire’s problem-posing approach to solicit community participation. With that goal, we next pose a set of seed problems for platform design. Problem-posing for information experience design should follow in parallel, but should be initiated by community members in context of their own relations and struggles, and is out of scope of our current work.
@@ 4. Problem-Posing for Emancipatory IA
The problems for platform design are chosen to prompt critical reflection, ideation, prototyping, experimentation, and critical studies. We intend for each of these problems to solicit a diverse range of sociotechnical design ideas that foregrounds our emancipatory aspirations. Each problem may be considered in isolation, but ultimately we seek proposals that can jointly address these concerns. We aim for this framework to expose a rich design space of alternative approaches that communities can explore to challenge the status quo. We pose our problems into two thematic groups, focusing on (1) platform operations and governance and (2) safeguards from capture and co-option.
4.1. Theme: Platform operation and governance
@@
The problems posed under this theme aim to provoke IA researchers and developers to rethink the design of our technological stack to create the conditions for community co-option and co-construction of IA technologies. Supporting community co-construction also motivates the need to explore democratic governance frameworks, effective policy enforcement mechanisms, and research and tooling support for downstream information experience design.
@@
    How do we design and develop the platform from the margins to legitimize community co-option and co-construction?
The platform should legitimize its own co-option and co-construction through strategic underdesigning. Community members should then be supported to build new IA artifacts (e.g., information retrieval components, moderation services, and information experiences). Communities should be able to leverage each other’s artifacts, and may also build competing artifacts if existing ones are inadequate or deemed harmful to community members. Through these interactions the platform should facilitate intra- and inter-community negotiations. The platform should also ensure an equal playing-field for community members; participation in co-construction should not be biased by who has access to more resources and labor (Neophytou and Mitra, 2026). Therefore, models of engaging communities should be informed by theories of power and strive for epistemic justice. Finally, to ensure the platform does not extract free labor from already marginalized communities, it should consider mechanisms for their compensation and long-term sustainability of this work.
@@
Community participation is not a miracle-cure for all societal concerns with IA. The opportunity to co-opt and co-construct may be leveraged by oppressors as well as the oppressed. We see these tensions already in context of app stores where their ownership by large for-profit corporations often imply that applications which harm marginalized communities are allowed to exist (Molla, 2019) while those that oppose authoritarian violence are pulled down under government pressure (Allyn, 2025). Addressing these with appropriate cultural and sociopolitical nuance requires robust democratic governance frameworks with community participation, which is distinct from participation for platform development.
All platforms need to moderate (Gillespie, 2018). Platform governance need to deal with critical moderation policies and mediate challenging cases, e.g., the case of the ‘Napalm girl’ photo (Levin et al., 2016). Content moderation and governance are not at odds with emancipatory aspirations; mitigating real-world harms is a core tenet of anti-oppression and emancipatory praxis. The challenge is to develop critical frameworks for democratic decision-making that are robust to capture.
@@
Democratic decision-making should be complemented by robust enforcement mechanisms. Our aspirations for community co-option and co-construction is not a loophole for circumventing moderation nor an excuse for harmful delays in mitigative actions—e.g., ensuring removal of child sexual abuse material (CSAM) from circulation. However, there is a natural tension here between developing infrastructure for the enforcement of policies and recreating infrastructure of censorship and centralized ideological control. To address this tension, we need innovations in sociotechnical infrastructure; and proposals for harm mitigation and justice that are restorative, reparative, and transformative (Kaba, 2021).
@@
    How do we support information experience design to promote emancipatory pedagogy?
Emancipatory IA challenges us to reimagine information experiences that are pedagogical and foreground critical scholarship. In the context of platform design, this motivates fundamental research and development of shared tools for communities to design information experiences in service of their own emancipation. Pedagogy in this context should not replicate the “banking concept of education”—nor act as systems for indoctrination—but encourage reflection, critical thinking, and dialog. This work should be informed by multidisciplinary perspectives—including HCI, design, critical theory, social and political sciences, humanities, and arts—and existing work in the area of search as learning (Urgo and Arguello, 2025).
4.2. Theme: Safeguards from capture and co-option
No liberatory struggle has the luxury to step outside of its own cage to architect a new liberated reality; the work begins from within the bounds of oppressive systems to transform the state of oppression to one of emancipation. The work on emancipatory IA should likewise begin situated within present-day structures of oppression. This motivates our work to prioritize safeguarding the platform from capture, blocking, political pressure, economic competition, and co-option.
@@ 4.2. Theme: Safeguards from capture and co-option
The platform should ensure that it cannot be pressured to take an action (or be blocked from taking an action) via state regulations or orders, through corporate or other institutional mandates, using economic or other incentives, by threats or acts of force or violence, or other forms of coercion. This requires that the infrastructure and its management, the governance of the platform, its sources of funding, and other aspects be decentralized to avoid exposing weak points that oppressive-actors can put pressure on to capture the platform. Any serious attempt at building an emancipatory IA platform should strategically consider adversarial forms of platform capture and develop necessary mitigation.
@@
The platform should support reasonably convenient mechanisms to circumvent attempts by governments or Internet service providers to block access to the service. There are precedents for online platforms—e.g., the messaging app Signal (Whittaker, 2024)—that provide mechanisms for circumventing authoritarian block on the app. Emancipatory IA should support similar mechanisms to ensure continued access to information under threats of authoritarian censorship.
@@
Any platform that succeeds in, or even attempts to, aid liberatory struggles would likely come under immediate political pressure. Beyond safeguarding against capture, the project of emancipatory IA should also consider how to ensure the physical and economic safety of those conducting the research and actively participating in development of the platform. This includes the safety of community members who participate in the co-construction of the platform. Safeguards may range from funding support and mutual aid, legal and regulatory protections, community norms, technological safeguards, and other mechanisms. Emancipatory IA platforms would also likely face severe economic competition from capitalist enterprises, and corporations might be particularly motivated to make the economics of the emancipatory IA platform untenable for its survival. Therefore, the economics of the platform should be a critical focus of the project.
@@
Emancipatory IA platforms risk being co-opted by adversarial actors. For-profit platforms, that do not share our emancipatory values, may attempt to extract economic value from the emancipatory IA platform without reciprocating. The platform design should be intentional in safeguarding against these risks and codify the need for reciprocity through technological design, and licensing and other legal means—e.g., Copyleft licensing (Frantsvog, 2012). Emancipatory IA should draw insights from other justice-oriented technology initiatives, such as the free software movement (Stallman, 1998).
@@ 5. The Salient Tensions of Emancipatory IA
Emancipatory IA situates itself within the tension between the oppressor and oppressed; and resolves that tension by making a normative commitment to side with the oppressed. However, attempting to simultaneously resolve several of the problems posed in Section 4 evokes other critical tensions—spanning the technological, social, political, economic, epistemic, legal, and personal. These tensions will manifest as we collectively put the theories to practice. In this section, we briefly discuss a short selection of such tensions that we need to negotiate.
Community participation vs. Agility
In the context of FD, there may be legitimate concerns about the feasibility of community participation at scale and its impact on the agility of platform development. While, such concerns may mirror known challenges in scaling PD (Zahlsen et al., 2022), it is crucial to recognize the structural differences with FD that may result in different outcomes. Unlike in PD, in FD there are no centralized group of technologists coordinating with communities. So, while it is difficult to predict how FD may scale with the number of participating communities without actually operationalizing FD, we speculate that it will depend on the structures of communication with and between communities facilitated by the platform design. Our argument is similar to that of Raymond’s in “The cathedral and the bazaar” (Raymond, 1999), where he counters Brooks’s law (“Adding manpower to a late software project makes it later”), arguing that the structure of coordination within the project determine how the complexity and communication cost grow with project members. Speculatively, the structure of coordination in FD may look as follows: Communities develop information experiences independently of each other; others may combine these different information experiences, without explicitly seeking permissions, to create richer compositions; while communities retain the right to “remove” their information experiences from specific compositions. Such distributed coordination may reflect “the bazaar” of information experiences, in contrast to today’s proprietary cathedrals of Google and Bing. If done right, this might scale better to meeting the needs of intersectional communities than the centralized design of proprietary platforms.
@@
In response to the growing concentration of power of a handful of platforms over public speech and IA, there is an emerging debate over the platform vs. protocol approach (Masnick, 2019). This is particularly dominant in the social media discourse where protocols like ActivityPub and ATProto have gained popularity as alternatives to billionaire-owned social media platforms. In this paper, we have liberally used the term “platform” to refer to the object of our envisioning in the context of emancipatory IA. This is intentional, partly to sidestep this debate earlier in the paper and use “platform” as a shorthand to refer to whatever solution that emerges as appropriate but also to reflect the realities of building for users who may not understand or care about the mechanics of platforms vs protocols, as we discuss below. While protocols present distinct advantages over platforms for user ownership of data, competition, and resistance to single points of capture, as Hendrix and Flowers (2022) argue, protocols in lieu of platforms do not inherently challenge the hierarchical social structures that they emerge from; the benefits that protocols do provide often rest on a neoliberal belief that market competition will drive better outcomes. However, Emancipatory IA nudges us to consider democratic governance designed explicitly to drive towards emancipatory outcomes.
@@
Platform governance (Gorwa, 2019) is a critical challenge for online platforms. Emancipatory IA not only inherits these challenges but should also be hardened against the risk of capture of any governance mechanism by authoritarian actors. It therefore requires additional safeguards and safety protocols, as well as contingency plans for when members of the governing body may be targeted or coerced. Policy enforcement mechanisms should navigate a similarly challenging landscape, balancing effectiveness of enforcement and the risks of implementing the tools for its own capture. This is one of the core tensions that future research on emancipatory IA needs to resolve.
@@
Design-related tensions emerge from opening up the platform to community co-option and co-construction. When large masses of Twitter users abandoned the social media site in 2022 to flock to Mastodon, they experienced significant growing pains onboarding onto the new platform (Diaz, 2022). New users had to shoulder the burden to familiarize themselves with a new platform that posed significant friction, especially for non-tech-savvy users, and exercise trust-based actions (such as selecting a Mastodon instance) that many felt ill-equipped to do (Stefanija, 2023). In enabling, co-option and co-construction, emancipatory IA should still ensure not to offload responsibilities to the user that might overwhelm them, and continue to prioritize the ease of usability of the platform. It should also avoid fragmentation and drive towards consistency of user experience. Fundamentally, it should drive a balance between the flexibility of design it affords to contributing communities and the overall usability of the platform.
@@
Lastly, we should be clear-eyed about challenges that emerge when the law diverges from justice. We opened this article with examples of authoritarian attempts at capture using executive orders and mandates. Under the conditions of oppression, the law disproportionately sides with the oppressor, which may create situations where the mandate of Emancipatory IA conflict with existing laws. Emancipatory IA platforms should find robust decision-making protocols to navigate these tensions while respecting legal frameworks (where possible), human rights, and our social commitment to do no harm.